OpenStudy (anonymous):

Solve this equation. g + 5 = 16 g = __

OpenStudy (anonymous):

G=11 11 + 5 = 16

OpenStudy (anonymous):

Heres the work: g + 5 - 5 = 16 - 5 g = 11

OpenStudy (anonymous):

g+5-5=16-5

OpenStudy (anonymous):

Already did but you are right

OpenStudy (anonymous):

Please don't give the answer, just help them get the answer. :)

OpenStudy (shadyykatie):

16-5

OpenStudy (anonymous):

Mason sorry to sound rude but theres no rule against giving answers and people usually come here for the answer not to be taught i will give how i got the answer but im not going to reteach them math or other subjects

OpenStudy (anonymous):

Please refer to the openstudy code of conduct~Give Help, Not Answers I will encourage and guide those needing help, and not just give them an answer

OpenStudy (anonymous):

Yes whatever the original equation was you do the opposite to both sides so: g + 5 = 16 would be g + 5 =5 = 16 - 5

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

^ no one cares

OpenStudy (anonymous):

If every moderator followed that there would be no open study users Mason get a Moderator to suspend or ban half their population for that go ahead

OpenStudy (anonymous):

Don't think they won't :) Because they will.

OpenStudy (anonymous):

make sure you explain it and see if the person gets the answer... if the person is really having trouble usually it's okay to give the answer, as long as you explain it correctly.

OpenStudy (anonymous):

@kewlgeek555 @Warriorz13

OpenStudy (anonymous):

Besides the pledge, we also have some specific guidelines on what not to do, outlined below. Failure to follow the guidelines may lead to your suspension or banning from the site. If you see any violations, please report them by using a report abuse link.

OpenStudy (warriorz13):

Well if that's the case then why does it say under the smart-score Questions Answered?

OpenStudy (anonymous):

@Mason_Jones573 we understand now ... you are not a moderator. Moderator's enforce the codes.

OpenStudy (warriorz13):

And also if that is the case then over %80 if the people that answer question of rhtis site have to get suspended or banned because that is all this site is for to get some last minute help

OpenStudy (anonymous):

Mason i have seen moderators give the straight answer at times so people really dont care considering moderators do it and you are the first to bring this up as i have seen so far

OpenStudy (anonymous):

I don't have to be a moderator to enforce the code.

OpenStudy (anonymous):

yes you do... you can report it, but the moderators enforce it.

OpenStudy (anonymous):

Then mason enforce it ban or suspend me right now go ahead.

OpenStudy (warriorz13):

But since you are not a moderator people will not take you seriously when you try to enforce the code!

OpenStudy (warriorz13):

And while your at it do the same to me!

OpenStudy (anonymous):

you report it to a moderator , not to yourself.

OpenStudy (anonymous):

You are not a moderator, in fact you are ranked lower then me @IsaiahCC100 and @Warriorz13 .

OpenStudy (warriorz13):

Then what is your rank?

OpenStudy (anonymous):

who?

OpenStudy (warriorz13):

Wait I am confused

OpenStudy (anonymous):

i'm 60 you're 62 and @Mason_Jones573 is 56

OpenStudy (anonymous):

Warrior he was talking to mason not us he just didnt use complete commas

OpenStudy (anonymous):

yes, I forgot the commas... lol

OpenStudy (warriorz13):

AND @IsaiahCC100 IS A 58 SO HE IS LOWER THEN YOU ARE And @IsaiahCC100 you don't have to keep telling me that since you sit right next to me

OpenStudy (warriorz13):

And plus the answer is what plus 5=16